i have to write an equation for the hyperbola that satisfies each set of conditions. if you could walk me thru one i think i can figure out the others. vertices (-4,0) and (4,0), conjugate axis length 8. Thanks
Answer by stanbon(75887)

5x^2+17x-12=0
-----------
Fartor:
---
5x^2 + 20x -3x -12 = 0
5x(x+4)-3(x+4) = 0
(x+4)(5x-3) = 0
--
x = -4 or x = 3/5
=======================